OpenStudy (anonymous):

The coordinates of the center of a circle with equation (x - 6)^2 + (y + 3)^2 = 7^2 are: A. (-6, 3) B. (-3, 6) C. (6, -3) D. (6, 7)

OpenStudy (anonymous):

A circle of center (a,b) and radius r has an equation (x-a)^2 + (y-b)^2 = r^2 Can you answer your question now?

OpenStudy (anonymous):

Compare the two formulae (x-a)^2 + (y-b)^2 = r^2 (x - 6)^2 + (y + 3)^2 = 7^2 What is a and b